# V-Max: A Reinforcement Learning Framework for Autonomous Driving

### v2

@@ 3.1 Rules of the Game
V-Max does not prescribe a universal goal for the policy; instead, it allows practitioners to define the desired behavior of the ego vehicle thanks to SDC (Self-Driving Car) paths. Waymax defines SDC path as the routes given to an agent by combining the logged future trajectory of the agent with all possible future routes after the logged trajectories. At the time of writing, these paths are not publicly available in WOMD. An alternative is to rely on expert-logged trajectories only as reference paths. However, this approach is problematic because expert trajectories represent privileged information that consistently demonstrates safe behavior, as shown in Figure 2. To overcome this limitation, V-Max enhances the simulation environment by incorporating reconstructed SDC paths. This addition enables researchers to define various practical tasks such as navigating to specific destinations or following predetermined routes.
@@
Figure 2: Illustration of potential limitations when utilizing expert trajectories as learning targets. The pink rectangle represents the ego, and the blue rectangles are the other vehicules. The blue path is one SDC path and the red cross is the last waypoint of the expert’s ground truth. Left: The trajectory terminates before a traffic signal, inadvertently encoding the implicit knowledge that the expert stopped at a red light. Center: The trajectory ends in the middle of the intersection, showing that the expert stopped to let other vehicles pass, which unintentionally teaches the policy when to yield. Right: The trajectory ends immediately prior to an intersection, which may result in the policy incidentally avoiding collisions by terminating at this location. In each scenario, we overlay the self-driving vehicle (SDC) path in blue, which provides a topologically consistent road representation without encoding such implicit behavioral biases, thus constituting a more appropriate supervisory signal for policy optimization.
@@ 3.2 Training RL Agents
Figure 3: Visualization of the observation process. Left: Scene-centered view of a scenario. SDC paths are displayed, in blue paths containing the expert trajectory, and in green showing alternative route options. Center: Ego-centric transformation with HD map filtering within a rectangular bounding box (70 meters front, 5 meters back, 20 meters on both sides of the SDC). Optionally, noise and masking can be applied to the perception of the scene. Right: Neural network input representation. Road boundaries are highlighted after roadgraph filtering. Only the eight closest objects to the SDC are retained, and the SDC path containing the ground truth is selected and interpolated into 10 points spaced 5 meters apart, providing a compact representation of the environment for decision making.
@@ 3.3 Evaluation and Benchmarking
The WOMD dataset does not provide speed limits, so to compute nuPlan’s speed-limit compliance metric, we inferred the speed limit from the expert trajectory using the following methodology. The dataset’s roads are located in San Francisco or Phoenix, where speed limits are one of $\{25,35,45,70\}$ mph. Additionally, road metadata indicates whether the agent is on a highway or in an urban scenario. We estimate the speed limit as 70 mph if and only if the agent is on a highway; otherwise, we assign the lowest speed limit such that the expert does not exceed it.
[…]
@@
To aggregate multiple metrics into a single score, we adopt the methodology from the nuPlan challenge Karnchanachari et al. (2024). Each episode is assigned a score based on a hybrid weighted average of all metric scores:
episode $\text{episode score}=\prod_{i\in\text{multiplier metrics}}\text{score}_{i}\times\sum_{j\in\text{average metrics}}\text{weight}_{j}\times\text{score}_{j}$
The complete list of metrics and their corresponding weights are provided in the supplementary material.
@@
Waymax includes a closed-loop reactive evaluation setup, where agents follow their logged trajectories but adjust their speed using an IDM policy Treiber et al. (2000).By default, all agents continue driving straight once they reach the end of their logged trajectory, regardless of road geometry. Additionally, stationary vehicles (e.g., those stopped at traffic lights or parked for the entire scenario) are initialized at the end of their logged trajectory. This causes them to start moving in a straight line, leading to unrealistic behaviors. These limitations make Waymax’s reactive evaluation setup unrealistic, so we chose not to include it in our experiments. nuPlan’s reactive agents also use an IDM policy, but use the roadgraph to generate their trajectories, resulting in more realistic behavior. We plan to integrate this feature in a future release.
@@
V-Max also integrates ReGentS Yin et al. (2024), a methodology for generating adversarial scenarios by modifying real-world driving data. In ReGentS, surrounding objects (e.g., vehicles, cyclists, and pedestrians) are optimized to create challenging situations for the agent while maintaining realistic and physically plausible interactions. The method prevents unrealistic swinging turns and unavoidable rear-end collisions, ensuring that the generated scenarios provide meaningful robustness evaluations.
@@ 4 Case Study: RL Design Choices for Autonomous Driving
We now present the practical functionalities of V-Max through distinct design choices for implementing an RL agent. To isolate the impact of each component, we use the control configuration defined in Table 1 and conduct experiments on observation functions, network encoders, reward shaping, and training datasets.
@@
Algorithm Observation Encoder Reward Training Dataset Evaluation Dataset
SAC Road LQ Navigation WOMD Training WOMD Valid
We replicate each run  [^1] Runs are executed on a single NVIDIA L4 GPU for 12-48 hours per run. across 3 random seeds, and report mean and standard deviation of evaluation metrics. Accuracy denotes the percentage of episodes completed without failure conditions (collisions, off-road, or traffic signal violations). Additionally, we report collision and off-road rates, along with Waymax’s progress ratio metric, which measures how far the agent progresses relative to the expert. Those metrics, as well as nuPlan’s metrics are aggregated in the V-Max score defined in Appendix A. Simulations are executed with a maximum of 64 objects per scenario to avoid memory overload.
@@
V-Max enables to configure the observation’s function with filters based on distance to the ego vehicle, and roadgraph types. We use this feature to benchmark 4 observation functions: (1) Base: includes all roadgraph points (2) Segment: selects closest road segments; (3) Lane: includes only lane centers; and (4) Road: includes only road boundaries.
All observation functions include the $N_{\mathrm{objects}=8}$ closest objects, $N_{\mathrm{lights}}=5$ closest traffic lights, and the same path target definition. This path target consists of a SDC path interpolated to 10 points spaced at 5-meter intervals. The four observations functions are illustrated in Figure 3.
Table 2: Performance of observation functions, evaluated using the control configuration (Table 1).
Observation Accuracy $\uparrow$ Collision $\downarrow$ Off-road $\downarrow$ Progress $\uparrow$ V-Max Score $\uparrow$
Base 96.92±0.20 1.94±0.18 0.91±0.09 173.32±4.21 0.85±0.00
Segment 96.15±0.35 1.80±0.29 1.83±0.22 152.85±15.12 0.84±0.01
Lane 95.99±0.42 2.18±0.41 1.60±0.06 136.67±10.79 0.84±0.00
Road 97.26±0.28 1.76±0.18 0.83±0.07 165.46±3.21 0.86±0.01
@@
We tested 3 rewards functions, starting with a basic reward that penalizes critical events (Safety), a reward that adds a signal to encourage following the target (Navigation), and third rewards that encourages to drive at the speed limit, maintain a safety distance to other vehicles and drive comfortably (Behavior).
@@
$r_{t}^{\mathrm{navigation}}$ $=r_{t}^{\mathrm{safety}}-0.6\cdot\mathbb{I}_{\mathrm{off-route}}(t)+0.2\cdot\mathbb{I}_{\mathrm{progress}(t)>\mathrm{progress}(t-1)}$
$r_{t}^{\mathrm{behavior}}$ $=r_{t}^{\mathrm{navigation}}+0.5\cdot\mathbb{I}_{\mathrm{comfort}}(t)+0.3\cdot\max\left(\frac{v(t)}{v_{\mathrm{lim}}},1\right)-0.6\cdot\mathbb{I}_{v(t)>v_{\mathrm{lim}}}$
Results in Table 4 show that the simple safety reward achieves high accuracy but results in a conservative policy, as reflected in its low progress score and overall V-Max score. Adding a navigation objective significantly improves progress while maintaining safety, leading to the highest overall performance. The behavior reward that adds an incentive to drive at higher speeds achieves the highest progress score, but at the cost of higher collision and off-road rates. This highlights the need to balance safety and efficiency when shaping a reward function.
Interestingly, while the behavior reward explicitly encourages comfort, it results in the lowest comfort score. This suggests that in our setting, the comfort signal is difficult to learn and is overshadowed by the stronger incentive to drive at higher speeds
@@ 5 Benchmark
We now present a benchmark of various approaches reimplemented in V-Max and report the best performance achieved by each method. Each method is executed across three random seeds, and the policy with the highest V-Max score is selected. The hyperparameters used for training are detailed in Appendix C.
@@ 5.1 Standard evaluation
We first compare approaches in the closed-loop non-reactive evaluation setup. PPO Schulman et al. (2017) and SAC Haarnoja et al. (2018) represent RL approaches, IDM Treiber et al. (2000) and PDM Dauner et al. (2023) represent rule-based approaches. Note that we used PDM-closed, as it has the best scores in closed-loop evaluation and is purely rule-based. To represent IL, we implemented BC with a Wayformer architecture, similar to the one used in Waymax’s benchmark. Finally, we also reimplemented BC-SAC from Lu et al. (2023).
@@
Table 6: V-Max benchmark, evaluations on WOMD Valid, in the closed-loop non-reactive setting. A complete definition of metrics is provided in Appendix A.
@@
Collision rate $\downarrow$ 0.56 7.88 5.95 1.74 2.16 7.81 13.14
Offroad $\downarrow$ 0.75 3.37 1.07 0.74 1.04 1.14 6.92
Red-light violation $\downarrow$ 0.61 0.24 0.21 0.09 0.19 0.33 0.55
Progress $\uparrow$ 97.30 140.44 138.12 169.01 159.61 189.52 86.87
At-fault collisions $\downarrow$ 0.49 3.10 1.00 1.17 1.46 5.57 4.49
Making progress $\uparrow$ 99.41 96.90 96.74 98.37 97.87 97.58 87.97
Driving direction compliance $\uparrow$ 97.91 99.46 99.55 98.75 97.90 99.26 97.47
TTC within bound $\uparrow$ 99.21 93.02 95.99 92.48 95.24 83.63 94.79
Progress along route ratio $\uparrow$ 99.09 91.99 90.84 95.49 93.32 94.61 78.64
Speed limit compliance $\uparrow$ 99.97 100 100 99.53 99.78 98.35 98.55
Multiple lanes score $\uparrow$ 91.26 99.08 96.79 93.00 93.00 94.00 89.68
Comfort $\uparrow$ 80.16 55.25 33.58 53.18 52.74 18.69 93.32
Accuracy $\uparrow$ 98.06 88.54 92.77 97.44 96.61 90.75 79.42
Accuracy only at-fault $\uparrow$ 98.14 93.32 97.72 98.00 97.32 88.07 92.99
V-Max score $\uparrow$ 0.93 0.85 0.85 0.88 0.86 0.78 0.72
@@
We see that V-Max enabled us to train a highly performing driving policy using a standard RL algorithm, achieving an accuracy of $97.44\%$ and strong performance across all sub-metrics of the V-Max score, using a single NVIDIA L4 GPU for 36 hours.
Notably, we see that our agents still fail to optimize for the comfort metric, except whose great comfort score can be explained by both the fact it learns "human-like" driving, but also because it learned a cautious policy, with poor progress scores.
5.2 Robustness assessment
In Table 7, we compare the impact of noisy initialization on SAC, BC, and PDM. As expected, BC struggles to recover from the perturbations, which push it into out-of-distribution states (covariate shift), leading to a significant drop in performance. In contrast, both RL (SAC) and the rule-based PDM only experience a small performance degradation, demonstrating greater robustness to initial perturbations.
Table 7: Performance of our best RL agent, BC and PDM in the noisy initialization setup. Scores on the first 294 scenarios of WOMD valid.
Algorithm Evaluation Accuracy $\uparrow$ Collision $\downarrow$ Offroad $\downarrow$ Progress $\uparrow$ V-Max score $\uparrow$
SAC Non-reactive 97.40 1.70 0.74 169.01 0.88
Noisy Init 94.50 3.00 1.70 162.29 0.83
BC Non-reactive 84.64 8.87 5.8 90.18 0.75
Noisy Init 35.15 32.76 27.64 52.65 0.25
PDM Non-reactive 93.50 4.40 1.70 152.17 0.82
Noisy Init 91.5 5.5 2.4 158 0.79
In Figure 4 we describe how our RL agent reacts to an adversarial agent generated using ReGentS Yin et al. (2024). To achieve a complete robustness assessment with this methodology, one should fist select a subset of the dataset it could be applied to. Indeed, in most scenarios (red lights, absence of close surrounding objects) the method is not applicable. We leave this to future work.
Refer to caption Refer to caption Refer to caption
(a) Scene-centered view of the initial scenario.
Refer to caption Refer to caption Refer to caption

### v3

@@ 2.1 Reinforcement Learning for Autonomous Driving
The successful application of RL in the simpler simulated racing task (Jaritz et al. (2018); Wurman et al. (2022)), further motivates the idea that performing simulators are key to unlock the potential of RL.
@@ 3.1 Rules of the Game
V-Max does not prescribe a universal goal for the policy; instead, it allows practitioners to define the desired behavior of the ego vehicle thanks to SDC (Self-Driving Car) paths. Waymax defines SDC path as the routes given to an agent by combining the logged future trajectory of the agent with all possible future routes after the logged trajectories. At the time of writing, these paths are not publicly available in WOMD. An alternative is to rely on expert-logged trajectories only as reference paths. However, this approach is problematic because expert trajectories represent privileged information that consistently demonstrates safe behavior, as shown in Figure 2. To overcome this limitation, V-Max enhances the simulation environment by incorporating reconstructed SDC paths. This addition enables researchers to define various tasks such as navigating to specific destinations or following predetermined routes.
@@
Figure 2: Illustration of potential limitations when using expert trajectories as learning targets. The pink rectangle represents the ego, and the blue rectangles are the other vehicules. The blue path is one SDC path and the red cross is the last waypoint of the expert’s ground truth. Left: The trajectory terminates before a traffic signal, inadvertently encoding the implicit knowledge that the expert stopped at a red light. Center: The trajectory ends in the middle of the intersection, showing that the expert stopped to let other vehicles pass, which unintentionally teaches the policy when to yield. Right: The trajectory ends immediately prior to an intersection, which may result in the policy incidentally avoiding collisions by terminating at this location. In each scenario, we overlay the self-driving vehicle (SDC) path in blue, which provides a topologically consistent road representation without encoding such implicit behavioral biases, thus constituting a more appropriate supervisory signal for policy optimization.
@@ 3.2 Training RL Agents
Figure 3: Visualization of the observation process. Left: Scene-centered view of a scenario. SDC paths are displayed, in blue paths containing the expert trajectory, and in green showing alternative route options. Center: Ego-centric transformation with HD map filtering within a rectangular bounding box (50 meters front, 5 meters back, 20 meters on both sides of the SDC). Optionally, noise and masking can be applied to the perception of the scene. Right: Neural network input representation. Road boundaries are highlighted after roadgraph filtering. Only the eight closest objects to the SDC are retained, and the SDC path containing the ground truth is selected and interpolated into 10 points spaced 5 meters apart, providing a compact representation of the environment for decision making.
@@ 3.3 Evaluation and Benchmarking
To aggregate multiple metrics into a single score, we adopt the methodology from the nuPlan challenge Karnchanachari et al. (2024). Each episode is assigned a score based on a hybrid weighted average of all metric scores. The complete list of metrics and their corresponding weights are provided in the supplementary material.
@@
In order to enable reactive evaluations, where the other vehicles are controlled by a policy rather than replaying their logged trajectory, we introduce a multi-agent version of WOMD. To construct it, we first define controllable vehicles as cars that are not offroad, and has data for the whole scenario (excluding cars that appear and disappear from the data because they were too far to be consistently detected). We then extract from the validation dataset (41872 scenarios) all scenarios with at least one controllable agent, and obtain a dataset of 31379 scenarios, with a mean of 8.72 controllable agents per scenario.
For non-ego agents, traffic lights data may be of poor quality, as the data was originally recorded from the ego vehicle, we robustify traffic lights by infering it from vehicles movement, similar to what was done in the nuPlan dataset Karnchanachari et al. (2024). Yan et al. (2025) went a step further by adding missing traffic lights in intersections in the WOMD dataset, which is planned to be integrated in V-Max.
Then, we construct SDC paths for the identified controllable agents, resulting in a new dataset, in which any policy developed on V-Max can be used to drive the controllable agents. It enables to perform closed-loop reactive evaluation with IDM policies similar to nuPlan and Waymax, excepted our IDM policies respect traffic lights, and follow the road rather than logged-trajectories.
This also enables us to try putting various policies like PDM or even evaluating how our RL policy drives when other agents use the same policy. We believe this could serve as a foundation for researchers trying to develop adversarial policies for robustness assessments.
@@
Finally, V-Max also integrates ReGentS Yin et al. (2024), a methodology for generating adversarial scenarios by modifying real-world driving data. In ReGentS, surrounding objects (e.g., vehicles, cyclists, and pedestrians) are optimized to create challenging situations for the agent while maintaining realistic and physically plausible interactions. The method prevents unrealistic swinging turns and unavoidable rear-end collisions, ensuring that the generated scenarios provide meaningful robustness evaluations.
@@ 4 Case Study: RL Design Choices for Autonomous Driving
We now present the functionalities of V-Max through distinct design choices for implementing an RL agent. To isolate the impact of each component, we use the control configuration defined in Table 1 and conduct experiments on observation functions, network encoders, reward shaping, and training datasets.
@@
Algorithm Encoder Reward Training Dataset Evaluation Dataset
SAC LQ Navigation WOMD Training WOMD Valid
We replicate each run [^1] Runs are executed on a single NVIDIA L4 GPU for 12-48 hours per run. across 3 random seeds, and report mean and standard deviation of evaluation metrics. Accuracy denotes the percentage of episodes completed without failure conditions (collisions, off-road, or traffic signal violations). Additionally, we report collision and off-road rates, along with Waymax’s progress ratio metric, which measures how far the agent progresses relative to the expert. Those metrics, as well as nuPlan’s metrics are aggregated in the V-Max score defined in Appendix B. Simulations are executed with a maximum of 64 objects per scenario to avoid memory overload.
@@
We perform an ablation study (Table 10 in Appendix) to understand which BEV features are most important for a RL agent.
We find that increasing temporal context from 1 to 5 observation steps does not always improves V-Max scores, and adding lane center or road line features does not systematically improve performance. It suggests that road edge with waypoints information alone provides sufficient spatial context for many scenarios in the context of the simulation. The best configuration (5 steps, 16 objects, road edge features, waypoints) achieves 0.88 V-Max score compared to 0.84-0.87 for other variants. All observation functions are illustrated in Figure 4.
@@
We developed a hierarchical reward structure consisting of three levels: Safety (basic collision and violation penalties), Navigation (adding progress incentives), and Behavior (incorporating comfort and speed compliance). To optimize reward weights, we conducted a grid search detailed in Table 12 in the Appendix.
@@
$r_{t}^{\mathrm{navigation}}$ $=r_{t}^{\mathrm{safety}}-0.2\cdot\mathbb{I}_{\mathrm{off-route}}(t)+0.2\cdot\mathbb{I}_{\mathrm{progress}(t)>\mathrm{progress}(t-1)}$
$r_{t}^{\mathrm{behavior}}$ $=r_{t}^{\mathrm{navigation}}+0.2\cdot\mathbb{I}_{\mathrm{comfort}}(t)-0.1\cdot\mathbb{I}_{v(t)>v_{\mathrm{lim}}}$
Results in Table 3 show that the simple safety reward achieves high accuracy but results in a conservative policy, as reflected in its low progress score and overall V-Max score. Adding a navigation objective significantly improves progress while maintaining safety. The behavior reward adds an incentive to drive with respects to the speed limits and the comfort of vehicle.
Reward shaping introduces significant optimization challenges due to competing objectives between reward components. Maximizing progress along the route while maintaining driving comfort creates a fundamental trade-off that requires precise speed control. High comfort weights can produce overly cautious policies with reduced speed, while high progression weights lead to aggressive driving that reduces safety. This demonstrates the challenge of multi-objective optimization in autonomous driving: individual reward terms can conflict with each other, which might be inefficient in order to achieve the desired driving behavior.
@@ 5 Benchmark
We now present a benchmark of various approaches reimplemented in V-Max and report the best performance achieved by each method. Each method is executed across three random seeds, and the policy with the highest V-Max score is selected. The hyperparameters used for training are detailed in Appendix E.
@@ 5.1 Standard evaluation
We first compare approaches in the closed-loop non-reactive evaluation setup. PPO Schulman et al. (2017) and SAC Haarnoja et al. (2018) represent RL approaches, IDM Treiber et al. (2000) and PDM Dauner et al. (2023) represent rule-based approaches, more details on our implementation of PDM can be found in Appendix C. Note that we used PDM-closed, as it has the best scores in closed-loop evaluation and is purely rule-based. To represent IL, we implemented BC with a Wayformer architecture, similar to the one used in Waymax’s benchmark. Finally, we also reimplemented BC-SAC from Lu et al. (2023), trained with the Safety reward.
@@
Table 5: V-Max benchmark - evaluations on WOMD Valid, in the closed-loop non-reactive setting. A complete definition of metrics is provided in Appendix B.
@@
Collision rate $\downarrow$ 0.49 2.78 0.96 0.89 1.25 7.81 13.14
Offroad $\downarrow$ 0.53 2.47 1.05 0.65 0.55 1.14 6.92
Red-light violation $\downarrow$ 0.65 0.26 0.21 0.21 0.20 0.33 0.55
Progress $\uparrow$ 98.94 144.44 140.58 170.85 90.98 189.52 86.87
At-fault collisions $\downarrow$ 0.49 2.78 0.96 0.89 1.25 5.57 4.49
Making progress $\uparrow$ 99.45 97.11 97.78 99.02 90.49 97.58 87.97
Driving direction compliance $\uparrow$ 97.82 99.66 99.52 98.11 90.95 99.26 97.47
TTC within bound $\uparrow$ 99.16 92.86 95.90 96.12 96.96 83.63 94.79
Progress along route ratio $\uparrow$ 99.18 92.65 92.17 95.13 75.60 94.61 78.64
Speed limit compliance $\uparrow$ 99.91 99.67 99.70 97.69 99.59 98.35 98.55
Multiple lanes score $\uparrow$ 90.96 99.25 96.28 90.62 72.65 94.00 89.68
Comfort $\uparrow$ 96.92 69.75 58.05 67.84 68.95 18.69 93.32
Accuracy $\uparrow$ 98.26 89.96 93.60 97.86 97.40 90.75 79.42
Accuracy only at-fault $\uparrow$ 98.33 94.51 97.78 98.25 98.02 88.07 92.99
V-Max score $\uparrow$ 0.945 0.876 0.887 0.892 0.713 0.784 0.721
@@
We see that V-Max enabled us to train a highly performing driving policy using a standard RL algorithm, achieving an accuracy of $97.86\%$ and strong performance across all sub-metrics of the V-Max score, using a single NVIDIA L4 GPU for 48 hours.
Although a RL planner can achieve expert-like accuracy in simulation, we see that our agents still fail to optimize for the comfort metric and the overall V-Max score, which could be investigate in future works in order to push a RL planner im the real world.
5.2 Reactive Evaluations
Table 6 presents our first experiment on using our multi-agent dataset for reactive evaluations. We see that safer models, like PDM and our SAC policy, allow to reduce collision rates for all policies. We specifically see how we managed to mitigate the weakness of log-replay by looking at the "Other collisions" raw, that account for collisions not imputable to the ego (rear-bumper collisions mainly).
Table 6: Reactive evaluations – Results on the first 1000 scenarios of our multi-agent dataset. Expert corresponds to the non-reactive setting.
Metric IDM vs PDM vs SAC vs
Expert IDM PDM SAC Expert IDM PDM SAC Expert IDM PDM SAC
Accuracy $\uparrow$ 89.7 91.0 96.0 95.5 94.3 95.1 98.4 97.8 97.7 95.9 98.5 98.4
At-Fault Collision $\downarrow$ 2.4 3.4 1.9 1.8 0.2 0.4 0.1 0.3 0.7 1.1 0.6 0.4
Other collisions $\downarrow$ 6.10 4.10 0.20 1.00 4.70 3.20 0.30 0.90 0.60 2.00 0.10 0.50
5.3 Adversarial Evaluation
Table 7 showcases how ReGentS can be used to perform adversarial evaluations. The ReGentS agents are optimized to have a trajectory that aggressively cuts-in the ego. We see that the RL agent, thanks to its closed-loop training, is far more robust than the other counterparts.
However, as it could be expected, most of the collisions created by ReGentS are classified as not imputable to the ego, but they should be taken into account in this setting, as such aggressive behaviors can occur in real life.
Table 7: Adversarial evaluations – Results on the first 100 valid scenario for ReGentS.
Metric IDM PDM SAC
Non Reactive ReGentS Non Reactive ReGentS Non Reactive ReGentS
Accuracy $\uparrow$ 92 43 96 53 98 76
At-Fault Collision $\downarrow$ 1 5 0 11 2 11
Other collisions $\downarrow$ 5 51 4 33 0 10